OpenStudy (anonymous):

What is the area of triangle ABC if a = 19, b = 14, and c = 23? in unites.

OpenStudy (anonymous):

I've had these questions multiple times, and I still don't know how to do it, so an explanation with it please.

OpenStudy (mathstudent55):

Have you heard of Heron's formula?

OpenStudy (anonymous):

I was assigned pre-cal without every knowing how to do it, so no.

OpenStudy (mathstudent55):

Heron's formula is a simple formula that gives you the area of a triangle given all its sides.

OpenStudy (mathstudent55):

For a triangle with side lengths a, b, and c, the area is \(A = \sqrt{s(s - a)(s - b)(s - c)} \) where \(s = \dfrac{a + b + c }{2} \)

OpenStudy (anonymous):

What is s ?

OpenStudy (mathstudent55):

Read the entire response above.

OpenStudy (anonymous):

Oh okay, I got it, thank you.

OpenStudy (anonymous):

Let me plug in, and I'll let you know what I got.

OpenStudy (anonymous):

132.82 is the answer.

OpenStudy (anonymous):

@mathstudent55 I have another question similar to this, but it's asking for degree, could you give me the formula for it?

OpenStudy (anonymous):

In triangle ABC, what is the measure of A if a = 9, b = 14, and c = 8?

OpenStudy (mathstudent55):

Correct. A = 132.8

OpenStudy (mathstudent55):

For the second question, since you have the measures of all sides but no angles, you need to use the law of cosines. \(a^2 = b^2 + c^2 - 2bc \cos A\)

OpenStudy (mathstudent55):

You know a, b, and c. The only unknown is A, which you can solve for using the inverse cosine function.

OpenStudy (anonymous):

Let me see if I can get this one, but I want double assurance.

OpenStudy (anonymous):

Wait how would I solve if I don't know what A is?

OpenStudy (mathstudent55):

First, replace all variables with their known values, leaving only A as A since you don't know it. Then solve for cos A.

OpenStudy (anonymous):

So I don't plug in cos A with the equation?

OpenStudy (anonymous):

9^2=14^2+8^2-2(14)(8) cos A

OpenStudy (mathstudent55):

\(a^2 = b^2 + c^2 - 2bc \cos A\) \(9^2 = 14^2 + 8^2 - 2\times 14 \times 8 \cos A\) You understand it so far?

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

Now do the squares, and multiply together the numbers that multiply the cos A.

OpenStudy (anonymous):

So I multiply cos A last?

OpenStudy (anonymous):

81=260-224 cos A

OpenStudy (mathstudent55):

Treat cos A as x, an unknown amount. \(81 = 196 + 64 - 224 \cos A\) \(81 = 260 - 224 \cos A\) Correct.

OpenStudy (mathstudent55):

We are solving for cos A. Now subtract 260 from both sides.

OpenStudy (anonymous):

224cos A=179?

OpenStudy (anonymous):

So 81=260-179.

OpenStudy (anonymous):

But that doesn't give me the degree.

OpenStudy (mathstudent55):

\(179 = 224 \cos A\) Now divide both sides by 224 \(\cos A = \dfrac{179}{224} \)

OpenStudy (mathstudent55):

We're getting there.

OpenStudy (mathstudent55):

Now we have solved for cos A, ok?

OpenStudy (anonymous):

0.799

OpenStudy (mathstudent55):

Now we use the inverse cosine function, which may be called \(\cos^{-1} \) in your calculator.

OpenStudy (anonymous):

So, Cos 0.799?

OpenStudy (anonymous):

Cos A= 0.799.

OpenStudy (mathstudent55):

Angle A has a cosine of 0.799. Now we want to know what angle A is. Do inverse cosine of 0.799.

OpenStudy (mathstudent55):

Your last line is correct. the previous one is not.

OpenStudy (mathstudent55):

You need to do \(\cos^{-1} 0.799 = \)

OpenStudy (anonymous):

cos^-1 0.799?

OpenStudy (anonymous):

0.645.

OpenStudy (anonymous):

Which is 36.97 degree.

OpenStudy (mathstudent55):

I don't know what you mean by 0.645, but A = 36.955 deg is correct.

OpenStudy (anonymous):

0.645 was result in radians.

OpenStudy (anonymous):

Thank you!

OpenStudy (mathstudent55):

Ok. You're welcome.

OpenStudy (mathstudent55):

BTW, if you need the other angles of this triangle, now you can use the law of sines to find the next angle. Then you can use the theorem that states that the sum of the measures of the angles of a triangle is 180 to find the last angle.

OpenStudy (anonymous):

Thank you, I'll refer back to this multiple times.

OpenStudy (mathstudent55):

You're welcome.